How can I write a linear system ( system of equations) corresponding to the matrix:
x y z
2 0 2 4
0 1 0 6
0 0 2 6

Answer by AnlytcPhil(1810)


How can I write a linear system ( system of equations) corresponding to the
matrix:
x y z
2 0 2 4
0 1 0 6
0 0 2 6
`
Just use the numbers below each letter as the coefficient of that letter,
indicate their addition, then put an equal sign, and then the number on the far
right
`
2x + 0y + 3z = 4
0x + 1y + 0z = 6
0x + 0y + 2z = 6
`
Now erase all the 0 terms and the 1 coefficient
`
2x ` ` `+ 3z = 4
` ` ` y ` ` `= 6
` ` ` ` ` 2z = 6
`
or just
`
2x + 3z = 4
y = 6
2z = 6
`
Edwin
AnlytcPhil@aol.com